OpenStudy (anonymous):

what is 0.28 as a fraction in simplest form

OpenStudy (chosenmatt):

the percent for 0.09 will be 9% so as a fraction 9 will be 9/100 so what will 0.28 be as a fraction?

OpenStudy (anonymous):

0.28=28/100=7/25